shift summary 13 June 1999, evening.

Nickolas Solomey (solomey@hep.uchicago.edu)
Mon, 14 Jun 1999 00:04:11 -0500

Shift Summary 13 June 1999
On shift: Patty McBride Nick Solomey
Others present: Rick**3, Ed. B., Sydney, Elizabeth, Bruce, DaveN and Peter

Hours of beam delivered at all intensities: 0.4
Hours of beam delivered at nominal intensity: 0
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 1 Hr laser scan.

Accounting of lost time: Tevatron recovery from a quench, several
times.
========================================================================
Usfull Runs ended this shift:
13571 A good laser scan
------------------------------------------------------------------------
Tapes brought to FCC this shift: None
========================================================================
Accesses during shift: Rick K. and Elizabeth
========================================================================
Overview of shift: Essentially No beam.
========================================================================
Description of Shift:
15:30> Arrived on shift, found Tevatron trying to recovery from a quench
at 15:00 hrs. Rick C. says he stills has a half hour of his primary
colimator position check studies to do after the beam comes back, then
we are clear to set up doing E832 data collection.

16:16> MCR reports quency recovery done, starting up, beam in 30 min.

16:40> Ed installed new CSQ6 (Q/E) constants from online gain calibartion.
He thinks that turning blower on and off may be disturbing fibers.

16:50> Gas check list done, He pressure low may have to open two more
tubes tomarrow.

17:15> Called MCR, it appears that even though the temperature has
recovered since the quench, something else has died during the quench with
a power supply and they are awaiting the arrival of an expert to tell them
what is wrong. They will post more information when they have an estimate.

18:07> Experts working on B2 dump SCR power supply. 45 min. estimate.

20:30> An error saying "The spill number does not change in 722 second"
during the starting of a run, this means that the slow daq was not
operating, but this was not true so Theo should look into it when he gets
a chance.

21:00> Dave N. opened one more tube of He. (Pressure = 120 psi.)

21:05> Beam starts to trickle in, Rick Ford paged.
21:25> Rick Ford starts his studies, comparing momentum spectra
of the two neutral beams.

21:29> Tevatron quenched at E1 and A3, estimate to come.

22:10> With the tevatron looking down and at most low intensity until they
get this quenching problems solved, Ed decided it was best to go to muon
mode. Step 1: zero magnet for muon run with an832negtopos ronald script on
epicure window, Step 2: called MCR and asked them to put our beam stop in,
Step 3: push our beamstop crash button, Step 4: Ask MCR to set up our
magnets for muon run. Now awaiting muons in muon mode.

23:05> Rick and Elizabeth made an access to fix the dead RC10B "extra"
trigger source. The discriminator width was too narrow and the anlalog out
did not reach 50 mV/hit. We adjusted all of the extra RC trigger source
widths to be about 1 bucket wide.

23:47> Beam returns, low intensity. Plan for the evening is to take muon
run, but Taku is on the phone with Sydney reprogramming the DAQ so it will
be a few minutes before the run is started. Paged Rick F. to let him know
beam is back since he said that if beam does not come back by midnight
with 2e12 or more he would not come in, but the beam came back before his
deadline with some minutes to spare.

23:59.999> Beam is back at 3.5e12! But it is muons for now.